### v3

# LLMs as Narcissistic Evaluators: When Ego Inflates Evaluation Scores

Abstract
Automatic evaluation of generated textual content presents an ongoing challenge within the field of NLP. Given the impressive capabilities of modern language models (LMs) across diverse NLP tasks, there is a growing trend to employ these models in creating innovative evaluation metrics for automated assessment of generation tasks. This paper investigates a pivotal question: Do language model-driven evaluation metrics inherently exhibit bias favoring texts generated by the same underlying language model? Specifically, we assess whether prominent LM-based evaluation metrics (e.g. BARTScore, T5Score, and GPTScore) demonstrate a favorable bias toward their respective underlying LMs in the context of summarization tasks. Our findings unveil a latent bias, particularly pronounced when such evaluation metrics are used in an reference-free manner without leveraging gold summaries. These results underscore that assessments provided by generative evaluation models can be influenced by factors beyond the inherent text quality, highlighting the necessity of developing more dependable evaluation protocols in the future.
LLMs as Narcissistic Evaluators: When Ego Inflates Evaluation Scores
[…]
1 Introduction
Evaluation is a fundamental element in both tracking progress and ensuring meaningful advancements across various dimensions within the field of Natural Language Processing (NLP). Therefore, the reliability of evaluation metrics plays a critical role in this process. Evaluating generated texts is one of the challenging and open problems in NLP given that different forms can convey the same meaning. This challenge has led to the development of various evaluation metrics for tasks involving Natural Language Generation (NLG). While human evaluation by experts stands as the most reliable approach for assessing generated outputs, it is costly and time-consuming, limiting its broader use. As a result, automatic evaluation metrics have emerged as practical alternatives to keep pace with the rapid progress in NLP van der Lee et al. (2019). Recent evaluation metrics for generation tasks, such as BERTScore Zhang et al. (2020), BARTScore Yuan et al. (2021), T5Score Qin et al. (2022), GPTScore Fu et al. (2023), and G-Eval Liu et al. (2023), increasingly rely on pretrained language models.
Figure 1: Examining the inherent bias within generative evaluation metrics towards outputs created by their underlying model reveals a clear existence of this bias. Our analysis shows that these metrics tend to assign inflated scores to outputs generated by the very model they are based on.
However, this trend poses a paradox, as the very outputs being evaluated are generated by these pretrained language models, raising concerns about inherent biases. For instance, an evaluation metric based on the BART model might yield inflated scores for outputs produced by a BART-based language model.
In this paper, we systematically investigate this potential bias, utilizing six prominent language models, namely BART Lewis et al. (2020), T5 Raffel et al. (2020), GPT-2 Radford et al. (2019), GPT-3 Brown et al. (2020), FLAN-T5 Chung et al. (2022), and Cohere along with their corresponding evaluation metrics (e.g. BARTScore, T5Score, and GPTScore) or conditional generative probability, for the task of summarization, which is a typical task in natural language generations and frequently employed in automatic text evaluation. Our analysis involved examining numerous variations of these six families of generative models, considering their varying sizes and finetuning settings both as generators and evaluators.
We conducted our analysis using the CNN/Daily Mail Hermann et al. (2015); Nallapati et al. (2016) and XSUM Narayan et al. (2018) summarization datasets. The assessment covers two settings: reference-based, using gold summaries for evaluation (a common approach in supervised summarization), and reference-free, comparing generated summaries against source documents (a common approach in both unsupervised summarization and factuality assessment).
Based on our analysis, we have derived the following findings: (1) Generative evaluators tend to assign higher scores to the content generated by the same underlying model. This bias becomes more pronounced when the fine-tuning configuration and model size match for both the generator and evaluator. (2) Inflated scores are particularly noticeable in the reference-free setting, which is concerning due to the popularity of this evaluation approach for assessing the factual correctness of generated texts (Koh et al., 2022). (3) Apart from the same model bias, inflated scores are also influenced by the preference for longer summaries by certain evaluators.
Our work has implications for model selection, evaluation strategies, and the development of more reliable and unbiased evaluation metrics in the field of natural language generation.
2 Related Work
Reference-based Evaluation Metrics
Reference-based metrics are commonly used to evaluate text generation tasks, including summarization, by measuring the similarity between generated and reference texts. Traditionally, metrics like BLEU Papineni et al. (2002) and ROUGE Lin (2004) were employed to assess a generated text based on surface-level similarities, measured through the n-gram overlap between the generated and reference texts. However, these traditional metrics often exhibit weak correlations with human judgments, particularly when evaluating advanced generation systems Peyrard (2019); Mathur et al. (2020).
Recognizing these limitations, recent trends in summarization evaluation lean towards semantic-level assessments, moving beyond direct word overlap comparisons. Notable metrics embracing this approach include BERTScore Zhang et al. (2020), MoverScore Zhao et al. (2019), BARTScore Yuan et al. (2021), BLEURT Sellam et al. (2020), and variations thereof. By leveraging pretrained language models, these metrics focus on capturing semantic content, providing a more nuanced and accurate evaluation of summarization system outputs.
Reference-free Evaluation Metrics
With the widespread use of generation models across diverse domains, the need for reference-free evaluation metrics has surged. In response to this challenge, recent attention has been directed towards metrics that enable the evaluation of generated texts solely based on source documents, especially when annotated reference texts may not be available for new domains Böhm et al. (2019); Gao et al. (2020); Chen et al. (2021); Wu et al. (2020); Scialom et al. (2021); Honovich et al. (2021); Zhong et al. (2022); Liu et al. (2023).
Representative reference-free metrics in recent years include generative evaluation models, exemplified by BARTScore Yuan et al. (2021) and GPTScore Fu et al. (2023), which are also used for reference-based evaluation. These metrics frame text evaluation as a natural language generation task, intuitively assigning higher probabilities to higher-quality generated texts. For instance, a recent study by Koh et al. (2022) has acknowledged BARTScore in reference-free mode as the factual consistency metric with the highest overall correlation to human factual consistency scores, particularly in the context of long document abstractive summarization. Therefore, the reliability of these metrics is important given their use for evaluating sensitive aspects such as factuality correctness.
Automatic Evaluation Metrics Pitfalls
Despite their widespread use, automatic evaluation metrics have notable shortcomings. These metrics may not be robust when faced with challenges such as spurious correlations, noise, or out-of-domain texts (Vu et al., 2022; Sai et al., 2021; He et al., 2023; Durmus et al., 2022). Furthermore, their effectiveness diminishes when evaluating very long documents (Amplayo et al., 2022). There is also evidence suggesting a potential bias towards ranking extractive summaries higher than abstractive ones Amplayo et al. (2022).
Traditional reference-based evaluation systems, such as ROUGE or BLEU, have been criticized for their inability to measure content quality or capture syntactic errors Reiter and Belz (2009). Consequently, these metrics may inaccurately reflect the reliability of generation systems. For example, they might assign high scores to outputs that are fluent but meaningless and unfaithful, as long as many of the same words are used Gehrmann et al. (2021). Although embedding-based metrics (e.g., BERTScore) show improved performance in similarity measurement, they are still inadequate for assessing the extent of shared information between two summaries, a crucial indicator of summary information quality Deutsch and Roth (2021).
Reference-free metrics, on the other hand, exhibit a bias towards outputs generated by models that are more similar to their own Deutsch et al. (2022). To the best of our knowledge, this study represents the initial attempt to perform a systematic exploration is yet to be undertaken. Additionally, question-answering-based reference-free metrics for summarization evaluation are prone to localizing errors within summaries Kamoi et al. (2023).
Metrics based on Large Language Models, which are capable of conducting both reference-based and reference-free evaluations, typically demonstrate superior correlations with human quality judgments across diverse NLG tasks and evaluation dimensions Deutsch et al. (2022). While prior work has reported that LLM-based metrics prefer LLM-generated text, raising a concern about the shortcomings of LLMs as evaluatiorsLiu et al. (2023), our work conducts a systematic evaluation to address a fundamental question: Do language model-driven evaluation metrics inherently display bias favoring texts generated by the same underlying language model? We explore this question across both reference-based and reference-free evaluations and for a range of different Large Language Models.
3 Methodology
To investigate the impact of the same model bias—determining whether a language model-based evaluator favors outputs generated by a similar language model—we conduct a comprehensive series of experiments involving both quantitative comparisons and qualitative analysis. Our quantitative comparisons involve using language models of varying sizes and finetuning configurations as both the evaluator and generator models. This structured approach enables us to systematically examine the potential bias across different LM configurations. Subsequently, we verify the results through qualitative analysis using a subset of models’ summaries that are accompanied by human evaluation to further demonstrate that higher scores produced by evaluators do not necessarily correlate with higher quality generated outputs.
3.1 Evaluators
We describe the generation progress as follows: given a source text s, and optional human written reference r , generate a hypothesis h, which can be represented as:
$y=f(\textbf{\emph{h}},a,\mathcal{S})$ (1)
where h denotes hypothesis, a refers to the aspect to evaluate, and $\mathcal{S}$ denotes supplementary text (i.e., s or r) that is employed alongside evaluations in various settings Fu et al. (2023). For instance, it could be source text s in a reference-free scenario which assesses the summary based on the source article directly Fabbri et al. (2021). Whereas in the reference-based paradigm, the evaluation considers semantic overlap between the generated and reference summaries r Bhandari et al. (2020).
The evaluators utilised in our study all share a conditional probability paradigm, which can be generally concluded as the definition in:
$\!\!\!\!\!\!\!\!\mathrm{Score}(\textbf{{h}}|d,a,\mathcal{S})=\sum_{t=1}^{m}w_{t}\log{p}(h_{t}|\textbf{{h}}_{<t},\mathcal{S},\theta),$ (2)
Under a sequence-to-sequence model parameterised by $\theta$, where $d$ refers to the task description and $w_{t}$ denotes the weight of the token at position $t$. Previous works normally treat each token equally Yuan et al. (2021); Fu et al. (2023); Finally, $\mathcal{S}$ is determined according to the aspect to evaluate.
BARTScore
BARTScore Yuan et al. (2021) introduced the generative evaluation approach treating text assessment as a generation task, employing probability of the text being generated by BART-based models Lewis et al. (2020) to assess the quality of text generated across various tasks such as machine translation, summarization, and data-to-text.
T5Score
T5Score Qin et al. (2022) was proposed providing both generative and discriminative training strategies for assessing T5-variant models as the core of this generative evaluation paradigm [^1] In our work, the training process of T5Score models only involves generative training due to the unavailability of publicly accessible checkpoints trained in a discriminative manner.. The integration of dual training strategies enables more types of data to be incorporated into the metric. T5Score closely aligns with BARTScore in terms of evaluation framework. Thus, when only considering the generative training strategy, T5Score is analogous to BARTScore, but for the T5 model series.
GPTScore
Leveraging generative models to conduct evaluation has been FURTHER ADVANCED with various of more recent large language models Fu et al. (2023), showing a great performance and covering a rich variety of aspects for comprehensive evaluations. With an understanding of natural language instructions, GPTScore (including GPT-X and FLAN-T5 models) can perform intricate and personalised assessments without requiring additional training.
Cohere
We additionally include Cohere, the more recent language model to enrich our experiment. The evaluation scores assigned by the model is calculated according to Eq. 2, aligned with BARTScore, T5Score, and GPTScore.
3.2 Generation Models
We analyze different variants of the BART, T5, GPT-2, GPT-3, FLAN-T5 and Cohere models, taking into account two different variables: the model size and the finetuning dataset. Regarding size, we consider small, base, medium, and large variations of each model, when available. For the finetuning dataset, we examine three distinct settings: (1) using the pretrained language model without finetuning on a summarization dataset, (2) finetuning on CNN, and (3) finetuning on XSUM. For instance, BART-Base-CNN represents a BART-base model that is finetuned on the CNN dataset. For each of the model types, we have used their corresponding standard prompts for the task of summarization. [^2] More details about the corresponding summarization prompts are included in Appendix A.2.
To ensure the reproducibility of our analysis, we exclusively employ publicly available checkpoints for the utilized models. Apart from the GPT3-Curie model that is taken from the OpenAI API [^3] Note that the newer models offered by the OpenAI API, such as GPT-4 , no longer support providing token probabilities, and hence they cannot be tested in our experiment. and generation model obtained from Cohere, the rest of the models are taken from the Hugging Face model hub [^4] https://huggingface.co/models..
We use each of these generation models both for generating the summaries [^5] We use the zero-shot setting for the models that are not finetuned on summarization datasets. as well as the underlying model for the LM-based evaluator. All the checkpoints used for generators and evaluators in our experiments can be found in the Appendix A (Table 4 and Table 5).
3.3 Datasets
We use documents from two well-established summarization datasets including CNN/DailyMail Hermann et al. (2015); Nallapati et al. (2016) and the extreme summarization (XSUM) dataset Narayan et al. (2018).
For quantitative comparisons, we randomly selected 500 documents from each of these datasets. We provide these documents to each of the generation models to obtain their corresponding generated summaries. For qualitative analysis, we use the SummEval benchmark Fabbri et al. (2021) and the RoSE benchmark Liu et al. (2022). These benchmarks include summaries from various generation models, as well as human evaluations, enabling us to assess the quality of these summaries.
The SummEval benchmark contains summaries generated by various summarization models for 100 articles from the CNN/DM test set, with each summary supplemented by human annotations. More specifically, SummEval incorporates human annotations by both expert and crowd-sourced human annotators, targeting dimensions of coherence, consistency, fluency, and relevance. Ratings are on a scale of 0 to 5, with higher values indicating better performance.
Similarly, RoSE contains summaries generated by recent generative models based on CNN/DM documents, accompanied by their corresponding human evaluations. We use 100 summaries from each of the BART and GPT-3 models from the ROSE benchmark. The RoSE benchmark proposed an assessment protocol termed ”Atomic Content Units” (ACUs). ACU score gauges quality of evaluated summaries based on whether the presence of single facts (i.e., atomic facts) from reference are included in the evaluated summaries. ACU score is calculated by ACU matching:
${f(s,\mathcal{A})=\frac{\lvert\mathcal{A}_{s}\rvert}{\lvert\mathcal{A}\rvert}}$ (3)
where $\mathcal{A}$ is a set of ACUs from gold summaries and s is the candidate summary.
Max Min Mean Median
RoSE-BART 1.00 0.00 0.37 0.38
RoSE-GPT3 0.90 0.00 0.27 0.25
SummEval-BART 5.00 2.67 4.56 4.67
SummEval-T5 5.00 2.33 4.46 4.67
SummEval-GPT2 5.00 1.33 3.52 3.33
Table 1: Distribution of human annotation scores on RoSE and SummEval datasets, where RoSE mainly considers the index ’ACU’ and SummEval focuses on ’Coherence’ and ’Consistency’ aspects according to expert annotators.
The distribution of human scores in RoSE and SummEval are given in Table 1.
Figure 2: Assessing Bias in the CNN/DM Dataset using heatmaps in the reference-free setting. Observing the darkest cells along the diagonal line, from the top left to the bottom right, indicates a distinct bias among evaluators towards their respective models. All evaluator scores are normalized to a range between 0 and 1. NB: the number in the bracket represents the average length of summaries (measured in words) produced by the respective model.
3.4 Quantitative Comparisons
We employ 20 language model-based evaluators for our experiments including six BARTScore evaluators Yuan et al. (2021), seven T5Score evaluators Qin et al. (2022), six GPTScore evaluators, and the Cohere evaluator. [^6] Appendix A.2 contains more details about the evaluators.
We assess the evaluators in two settings: (a) reference-free, where the metric evaluates the likelihood of the summary being generated from the source text, and (b) reference-based, where the generated summary is evaluated based on the reference summary.
Due to the nature of log probabilities, original scores from each evaluator should be negative, and a higher score indicates better quality according to the evaluator. When weights $w_{t}$ in Eq. 2 are treated equally, the evaluation protocols of BARTScore, T5Score, and GPTScore are all conditional probability paradigms. To ensure comparability among the scores provided by 20 distinguished evaluators, a uniform normalization process is applied to the scores generated by each evaluator. The normalization procedure standardizes the scores across a scale ranging from 0 to $\alpha$  [^7] In our work, we set parameter $\alpha$ to 1. as formulated in Eq. 4, where $S_{i,j}$ indicates scores evaluated by the $j$-th evaluator on summaries generated by the $i$-th generator.
$S_{i,j}^{norm}=\frac{\alpha(S_{i,j}-\min_{i}S_{i,j})}{\max_{i}S_{i,j}-\min_{i}S_{i,j}}$ (4)
In this context, a normalized score of $\alpha$ signifies the highest quality attributed by the generator, while a score of 0 indicates the lowest quality.
3.5 Qualitative Analysis
For qualitative analysis, we employ Spearman Correlations Zar (2014) and Kendall Correlations Freedman et al. (2007), which respectively assess monotonic relationships and order associations between human evaluations and LM evaluator scores.
For the SummEval dataset, we calculate the average value of expert annotations for coherence and consistency aspects, aligned with the reference-free input setting in the evaluation protocol specified in Yuan et al. (2021). For the evaluations based on the RoSE benchmark, ACU annotations are suited for reference-based summary salience evaluation. Therefore, we employ the correlation values obtained from the SummEval dataset for the reference-free setting and those from the RoSE benchmark for the reference-based setting.
4 Experimental Results
4.1 Quantitative Comparisons: Assessing Bias in LM-Evaluators Towards Their Own Output
Figure 3: Assessing Bias on CNN/DM Dataset using heatmaps in the reference-based setting. Observing darker cells along the diagonal line indicates potential bias. All evaluator scores are normalized to a range between 0 and 1. NB: the number in the bracket represents the average length of summaries (measured in words) produced by the respective model.
Figures 2 and 3 display heatmaps presenting evaluator scores for various summaries generated by different generators from CNN/DM documents. These scores are computed by averaging the individual scores of the selected 500 documents. In both heatmaps, we observe darker cells along the diagonal line, running from the top left to the bottom right. This indicates potential evaluator bias towards their corresponding generator models. However, this bias is notably more pronounced in the reference-free setting, commonly used for factuality evaluation Koh et al. (2022). In this context, we highlight the presence of the darkest cells along the diagonal line.
Furthermore, as shown in Figure 2, we note a distinct trend: T5-based generators, whether fine-tuned or not, tend to receive higher scores when assessed using different T5Score variations compared to evaluations using BARTScore, GPTScore, or Cohere. This results in a concentrated dark rectangle at the heatmap’s centre. Similarly, we observe a parallel trend for BART-based generators, whether fine-tuned or not. It is also quite evident that when using FLAN-T5 variants and Cohere as evaluators, they tend to favor the summaries generated by their respective models. Meanwhile, evaluators are prone to assigning higher ranks to generators under the same dataset compared to those fine-tuned on others. For instance, when utilizing T5 fine-tuned on the XSUM dataset as evaluators, a preference is observed for BART-XSUM generators over T5-vanilla models. We observe the same pattern on summaries generated based XSUM documents. The corresponding heatmap is included in Figure 6 in Appendix B.1.
It can also be observed that BART-based evaluators tend to rank FLAN-T5 generators the lowest, resulting in a normalized score of 0. This is likely to be attributed to the difference between BART and FLAN-T5’s architectures and BART-based models’ preference towards longer summaries. Similar observations can be observed that T5-based models score the summaries generated by Cohere low, as they tend to favor shorter summaries. Detailed analysis of the correlation between the length of generated summaries and the scores assigned by each evaluator is given in Figure 5.
Figure 4: Heatmaps of evaluation scores on the RoSE & SummEval benchmarks for the reference-free and reference-based setting. It is recommended to use the reference-free setting for SummEval and the reference-based setting for RoSE, aligning with the specific aspects each benchmark emphasizes.
Figure 5: Spearman Correlation between the length of generated summaries and the reference-free scores assigned by each evaluator. A higher positive score indicates that an evaluator prefers longer summaries, while a lower negative score indicates a preference for shorter summaries.
4.2 Bias towards longer summaries
Another notable pattern in Figure 2 is the high scores for the BART-based generators, indicated by both BARTScore variants and different GPTScores. To further investigate this phenomenon, we calculate the average length of summaries generated by each generator for each of the datasets. Notably, BART models and the Cohere that have not been fine-tuned for summarization tend to produce the longest summaries on average. This is followed by the fine-tuned BART models on the CNN dataset. Conversely, models fine-tuned on XSUM generate the shortest summaries. [^8] Note that the XSUM dataset contains one-sentence summaries.
Subsequently, we computed the Spearman correlation between the scores under the reference-free setting given by each of our examined evaluators and the length of the corresponding summary. The results are presented in Figure 5. Based on these results, with the exception of evaluators fine-tuned on XSUM, BARTScore and GPTScore variants tend to assign higher scores to longer summaries. This observation explains the darker squares positioned in the top-right corner of Figure 2 for high values of GPTScore variants, highlighting their inclination to assign higher scores to BART and BART-CNN generators that produce longer summaries. It is worth noting that this correlation with summary length is prominent within the reference-free setting. We observe a similar but less obvious pattern in the reference-based evaluations, as shown in Figure 3.
4.3 Qualitative Analysis: Correlation of Bias with Human Evaluation
To further verify the same model bias, we repeat the experiments from § 4.1 on summarization benchmarks that are accompanied by human evaluations. While the number of summaries in these benchmarks is limited compared to those in § 4.1, we can use the human annotations to verify that the inflated scores are not correlated with human evaluations.
Figure 4 shows the evaluation results for the SummEval and RoSE benchmarks for the reference-free and reference-based setting respectively. It is advisable to utilize the reference-free setting for SummEval and the reference-based setting for RoSE, aligning with the specific aspects each benchmark emphasizes Yuan et al. (2021). Overall, we observe a trend similar to that shown in Figure 2. For instance, the T5-base generator receives higher scores from T5-based evaluators. [^9] In SummEval, the T5 model is only ranked higher when evaluated with certain variants of the T5Score in the reference-based setting. Meanwhile, BART-based models receive higher scores from both BARTScore and GPTScore evaluators, instead of T5 evaluator.
For human correlations, we use the RoSE and SummEval benchmark for evaluating the correlation in the reference-based and reference-free setting respectively. Table 3 (in Appendix) presents the Spearman and Kendall correlation values of SummEval in the reference-free setting. Overall, we observe that none of the evaluators have a strong correlation with the human annotations on either of these benchmarks. Due to the limited size of the samples (i.e., 100 summaries from SummEval and 100 summaries from ROSE with human annotations, as described in §3.3) and the absence of many of our investigated generators in § 4.1, we cannot draw a conclusive conclusion from the correlation values. Nevertheless, these results demonstrate that none of these evaluators highly correlate with human annotations, and as observed in § 4.1, their inflated scores for their own underlying generator may contribute to this low correlation.
Evaluator RoSE - Reference-based
Spearman Kendall
BART-Base 0.454 0.310
BART-Large 0.298 0.218
BART-Base-CNN 0.488 0.345
BART-Large-CNN 0.468 0.329
BART-Base-XSUM 0.150 0.103
BART-Large-XSUM 0.371 0.253
T5-Small 0.396 0.284
T5-Base 0.395 0.285
T5-Large 0.392 0.282
T5-Small-CNN 0.393 0.281
T5-Base-CNN 0.391 0.276
T5-Small-XSUM 0.379 0.269
T5-Large-XSUM 0.462 0.324
GPT2 0.375 0.255
GPT2-Medium 0.357 0.244
GPT2-Large 0.353 0.242
GPT3-Curie 0.310 0.214
FLANT5-Base 0.460 0.325
FLANT5-XL 0.433 0.304
Cohere-Command 0.384 0.267
Table 2: Spearman and Kendall correlations between reference-based evaluation scores and human annotations using annotations in RoSE. The highest coefficient in each human annotation setting is bolded.
5 Conclusions
Based on experiments, we make the following conclusions:
First, the popularity of generative evaluation metrics, such as BARTScore, is on the rise for evaluating the factual accuracy of generated content—a critical concern in modern generator models. However, our results reveal that this evaluation approach is susceptible to the same model bias, highlighting the need for more robust metrics to assess factual correctness reliably. Second, our analysis indicates that models fine-tuned on the XSUM dataset are not suitable for direct integration into evaluators due to their bias towards shorter summaries. The exception is their use for evaluating summaries aligned with XSUM-style content. Third, notably, similar to traditional evaluation metrics Sun et al. (2019), contemporary evaluation metrics might also lean towards favoring longer summaries. This bias should be considered when interpreting and applying these metrics. Finally, our study uncovers the presence of the same model bias across all assessed evaluators. Consequently, we recommend avoiding the use of the same underlying model as the generator for assessment. Although the limited human evaluations for our examined models prevent definitive conclusions on selecting the best generative evaluator, our research charts a promising direction for designing more resilient and unbiased evaluation metrics.
In summary, our study identifies a new type of bias in generative evaluators encouraging future research in this direction for designing fairer evaluation metrics.
Limitations
We noted that our work has the following limitations. Firstly, our experiment has been focused on the summarization task. Expanding the evaluation to encompass a broader range of generation tasks would be highly beneficial. Secondly, conducting a larger-scale human evaluation would be advantageous, as our current experiments are constrained by the limited sample sizes from SummEval and ROSE. Finally, incorporating additional generation models and evaluators in future work would further enrich the experiment.
Ethics Statement
This paper raises no ethical concerns. The data and additional resources used in this study are open-source and widely employed in existing works.
Appendix A Evaluation Setting
Evaluator SummEval - Reference-free
Spearman Kendall
BART-Base 0.023 0.010
BART-Large 0.120 0.082
BART-Base-CNN 0.240 0.164
BART-Large-CNN 0.227 0.149
BART-Base-XSUM 0.081 0.055
BART-Large-XSUM 0.061 0.043
T5-Small 0.396 0.275
T5-Base 0.365 0.249
T5-Large 0.337 0.236
T5-Small-CNN 0.448 0.315
T5-Base-CNN 0.259 0.172
T5-Small-XSUM 0.235 0.167
T5-Large-XSUM 0.110 0.081
GPT2 0.153 0.107
GPT2-Medium 0.207 0.146
GPT2-Large 0.183 0.128
GPT3-Curie 0.326 0.228
FLANT5-Base 0.361 0.252
FLANT5-XL 0.433 0.305
Cohere-Command 0.328 0.227
Table 3: Spearman and Kendall correlations between reference-free evaluation scores and human annotations using annotations in SummEval. The highest coefficient in each human annotation setting is bolded.
A.1 Generator
Further information about models utilized in our generation models including checkpoint names, prompt setting, etc., are presented in Table 4.
Name of Generator Name of Checkpoint or Model Suffix Prefix
BART-Base facebook/bart-base Summarize:
BART-Large facebook/bart-large Summarize:
BART-Base-CNN ainize/bart-base-cnn
BART-Large-CNN facebook/bart-large-cnn
BART-Base-XSUM morenolq/bart-base-xsum
BART-Large-XSUM facebook/bart-large-xsum
T5-Small t5-small Summarize:
T5-Base t5-base Summarize:
T5-Large t5-large Summarize:
T5-Small-CNN ubikpt/t5-small-finetuned-cnn
T5-Base-CNN flax-community/t5-base-cnn-dm
T5-Small-XSUM pki/t5-small-finetuned xsum
T5-Large-XSUM sysresearch101/t5-large-finetuned-xsum
GPT2 openai-community/gpt2 TL;DR:
GPT2-Medium openai-community/gpt2-medium TL;DR:
GPT2-Large openai-community/gpt2-large TL;DR:
GPT3-Curie text-curie-001 TL;DR:
FLANT5-Base google/flan-t5-base TL;DR:
FLANT5-XL google/flan-t5-xl TL;DR:
Cohere-Command api.cohere.ai/v1/generate Write a concise summarization:
Table 4: Name of checkpoints or model utilized in our generation setting with corresponding prompt configurations, ’text-curie-001’ is the model name provided by OpenAI API, and ’api.cohere.ai/v1/generate’ denotes model names provided by Cohere API, alongside other checkpoints available through Hugging Face.
A.2 Evaluator
Further information about models utilized in our evaluation models are presented in Table 5.
Name of Evaluator Name of Checkpoint or Model Suffix Prefix
BART-Base facebook/bart-base Summarize:
BART-Large facebook/bart-large Summarize:
BART-Base-CNN ainize/bart-base-cnn
BART-Large-CNN facebook/bart-large-cnn
BART-Base-XSUM morenolq/bart-base-xsum
BART-Large-XSUM facebook/bart-large-xsum
T5-Small t5-small Summarize:
T5-Base t5-base Summarize:
T5-Large t5-large Summarize:
T5-Small-CNN ubikpt/t5-small-finetuned-cnn
T5-Base-CNN flax-community/t5-base-cnn-dm
T5-Small-XSUM pki/t5-small-finetuned xsum
T5-Large-XSUM sysresearch101/t5-large-finetuned-xsum
GPT2 openai-community/gpt2 TL;DR:
GPT2-Medium openai-community/gpt2-medium TL;DR:
GPT2-Large openai-community/gpt2-large TL;DR:
GPT3-Curie text-curie-001 TL;DR:
FLANT5-Base google/flan-t5-base TL;DR:
FLANT5-XL google/flan-t5-xl TL;DR:
Cohere-Command api.cohere.ai/v1/generate Write a concise summarization:
Table 5: Name of checkpoints or model utilized in our evaluation study with corresponding prompt configurations, ’text-curie-001’ is the model name provided by OpenAI API, and ’api.cohere.ai/v1/generate’ denotes model names provided by Cohere API, alongside other checkpoints available through Hugging Face.
Appendix B Evaluation Results
B.1 Reference-free Setting
Results of XSUM Dataset in Reference-free setting are presented in Figure 6. Evaluation scores for RoSE and SummEval benchmarks under the reference-free setting are shown in Figure 4.
For the meta evaluation, Spearman and Kendall correlation values in the reference-free setting for SummEval benchmark are shown in Table 3.
Figure 6: Assessing Bias in the XSUM Dataset using heatmaps in the reference-free setting. Observing the darkest cells along the diagonal line, from the top left to the bottom right, indicates a distinct bias among evaluators towards their respective models. All evaluator scores are normalized to a range between 0 and 1. NB: the number in the bracket represents the average length of summaries (measured in words) produced by the respective model.
B.2 Reference-based Setting
Heatmap of evaluation result on CNN/DM dataset under reference-based setting is given by Figure 3
Evaluation scores for RoSE and SummEval benchmarks under the reference-based setting are illustrated by Figure 4.
For the meta evaluation, Spearman and Kendall correlation values in the reference-based setting for RoSE benchmark are shown in Table 2.

### v4

@@
 BARTScore, T5Score, and GPTScore) demonstrate a favorable bias toward their respective underlying LMs in the context of summarization tasks.
-Our findings unveil a latent bias, particularly pronounced when such evaluation metrics are used in an reference-free manner without leveraging gold summaries.
-These results underscore that assessments provided by generative evaluation models can be influenced by factors beyond the inherent text quality, highlighting the necessity of developing more dependable evaluation protocols in the future.
+Our findings unveil a latent bias, particularly pronounced when such evaluation metrics are used in a reference-free manner without leveraging gold summaries.
+These results underscore that assessments provided by generative evaluation models can be influenced by factors beyond the inherent text quality, highlighting the necessity of developing more reliable evaluation protocols in the future.
 LLMs as Narcissistic Evaluators: When Ego Inflates Evaluation Scores
-[…]
+Yiqi Liu1 Nafise Sadat Moosavi2 Chenghua Lin1 1University of Manchester 2University of Sheffield yiqi.liu-6@postgrad.manchester.ac.uk n.s.moosavi@sheffield.ac.uk chenghua.lin@manchester.ac.uk
 1 Introduction
-Evaluation is a fundamental element in both tracking progress and ensuring meaningful advancements across various dimensions within the field of Natural Language Processing (NLP).
+Evaluation is a fundamental element in both tracking progress and ensuring meaningful advancements across various dimensions within the field of Natural Language Processing.
 Therefore, the reliability of evaluation metrics plays a critical role in this process.
@@
 We conducted our analysis using the CNN/Daily Mail Hermann et al.
-(2015); Nallapati et al.
 (2015) and XSUM Narayan et al.
@@
 (2) Inflated scores are particularly noticeable in the reference-free setting, which is concerning due to the popularity of this evaluation approach for assessing the factual correctness of generated texts (Koh et al., 2022).
-(3) Apart from the same model bias, inflated scores are also influenced by the preference for longer summaries by certain evaluators.
+(3) Apart from self-bias, inflated scores are also influenced by the preference for longer summaries by certain evaluators.
 Our work has implications for model selection, evaluation strategies, and the development of more reliable and unbiased evaluation metrics in the field of natural language generation.
@@
 (2002) and ROUGE Lin (2004) were employed to assess a generated text based on surface-level similarities, measured through the n-gram overlap between the generated and reference texts.
-However, these traditional metrics often exhibit weak correlations with human judgments, particularly when evaluating advanced generation systems Peyrard (2019); Mathur et al.
-(2020).
-Recognizing these limitations, recent trends in summarization evaluation lean towards semantic-level assessments, moving beyond direct word overlap comparisons.
+Recent trends in summarization evaluation lean towards semantic-level assessments, moving beyond direct word overlap comparisons.
 Notable metrics embracing this approach include BERTScore Zhang et al.
@@
 (2019); Gao et al.
+(2020); Wu et al.
 (2020); Chen et al.
-(2021); Wu et al.
 (2021); Scialom et al.
@@
 Despite their widespread use, automatic evaluation metrics have notable shortcomings.
-These metrics may not be robust when faced with challenges such as spurious correlations, noise, or out-of-domain texts (Vu et al., 2022; Sai et al., 2021; He et al., 2023; Durmus et al., 2022).
+These metrics may not be robust when faced with challenges such as spurious correlations, noise, or out-of-domain texts (Sai et al., 2021; Vu et al., 2022; Durmus et al., 2022; Zhao et al., 2023; He et al., 2023).
 Furthermore, their effectiveness diminishes when evaluating very long documents (Amplayo et al., 2022).
@@
 (2022).
-Traditional reference-based evaluation systems, such as ROUGE or BLEU, have been criticized for their inability to measure content quality or capture syntactic errors Reiter and Belz (2009).
-Consequently, these metrics may inaccurately reflect the reliability of generation systems.
+Traditional reference-based evaluation metrics such as ROUGE or BLEU have been criticized for their inability to measure content quality or capture syntactic errors Reiter and Belz (2009).
+Consequently, these traditional metrics often exhibit weak correlations with human judgements, demonstrating that they cannot accurately reflect the real-world performance of generation systems Peyrard (2019); Mathur et al.
+(2020).
 For example, they might assign high scores to outputs that are fluent but meaningless and unfaithful, as long as many of the same words are used Gehrmann et al.
@@
 (2022).
-To the best of our knowledge, this study represents the initial attempt to perform a systematic exploration is yet to be undertaken.
-Additionally, question-answering-based reference-free metrics for summarization evaluation are prone to localizing errors within summaries Kamoi et al.
+To the best of our knowledge, this study represents the initial attempt to perform an exploration, which has not yet been undertaken systematically.
+Additionally, question-answering-based reference-free metrics for summarization evaluation are prone to inheriting errors within summaries Kamoi et al.
 (2023).
-Metrics based on Large Language Models, which are capable of conducting both reference-based and reference-free evaluations, typically demonstrate superior correlations with human quality judgments across diverse NLG tasks and evaluation dimensions Deutsch et al.
+Metrics based on Large Language Models, which are capable of conducting both reference-based and reference-free evaluations, typically demonstrate superior correlations with human quality judgements across diverse NLG tasks and evaluation dimensions Deutsch et al.
 (2022).
-While prior work has reported that LLM-based metrics prefer LLM-generated text, raising a concern about the shortcomings of LLMs as evaluatiorsLiu et al.
-(2023), our work conducts a systematic evaluation to address a fundamental question: Do language model-driven evaluation metrics inherently display bias favoring texts generated by the same underlying language model?
+While prior work has reported that LLM-based metrics prefer LLM-generated text, raising a concern about the shortcomings of LLMs as evaluators Liu et al.
+(2023), our work conducts a systematic evaluation to address a fundamental question: Do language model-driven evaluation metrics inherently display bias favouring texts generated by the same underlying language model?
 We explore this question across both reference-based and reference-free evaluations and for a range of different large language models.
 3 Methodology
-To investigate the impact of the same model bias—determining whether a language model-based evaluator favors outputs generated by a similar language model—we conduct a comprehensive series of experiments involving both quantitative comparisons and qualitative analysis.
+To investigate the impact of the model’s self-bias—determining whether a language model-based evaluator favours outputs generated by a similar language model—we conduct a comprehensive series of experiments involving both quantitative comparisons and qualitative analysis.
 Our quantitative comparisons involve using language models of varying sizes and finetuning configurations as both the evaluator and generator models.
 This structured approach enables us to systematically examine the potential bias across different LM configurations.
-Subsequently, we verify the results through qualitative analysis using a subset of models’ summaries that are accompanied by human evaluation to further demonstrate that higher scores produced by evaluators do not necessarily correlate with higher quality generated outputs.
+Subsequently, we verify the results through qualitative analysis using a subset of models’ summaries that are accompanied by human evaluation to further demonstrate that higher scores produced by evaluators as a result of self-bias do not necessarily correlate with higher quality generated outputs.
 3.1 Evaluators
-We describe the generation progress as follows: given a source text s, and optional human written reference r , generate a hypothesis h, which can be represented as:
+We describe the evaluation process as follows: given a source text s, a human written reference r, generate a hypothesis h, which can be represented as:
 $y=f(\textbf{\emph{h}},a,\mathcal{S})$ (1)
@@
 (2023).
-For instance, it could be source text s in a reference-free scenario which assesses the summary based on the source article directly Fabbri et al.
+For instance, it could be the source text s in a reference-free scenario which assesses the summary based on the source article directly Fabbri et al.
 (2021).
-Whereas in the reference-based paradigm, the evaluation considers semantic overlap between the generated and reference summaries r Bhandari et al.
+Whereas in the reference-based paradigm, the evaluation considers semantic overlap between the generated hypothesis h (e.g.
+model generated summaries) and reference summaries r Bhandari et al.
 (2020).
-The evaluators utilised in our study all share a conditional probability paradigm, which can be generally concluded as the definition in:
-$\!\!\!\!\!\!\!\!\mathrm{Score}(\textbf{{h}}|d,a,\mathcal{S})=\sum_{t=1}^{m}w_{t}\log{p}(h_{t}|\textbf{{h}}_{<t},\mathcal{S},\theta),$ (2)
-Under a sequence-to-sequence model parameterised by $\theta$, where $d$ refers to the task description and $w_{t}$ denotes the weight of the token at position $t$.
-Previous works normally treat each token equally Yuan et al.
+The evaluators (i.e. based on BART, T5, GPT model variants as well as Cohere) utilised in our study all share a conditional probability paradigm, which can generally be formulated as
+$\!\!\!\!\!\!\!\!\mathrm{Score}(\textbf{{h}}|d,a,\mathcal{S})=\sum_{t=1}^{m}w_{t}\log{p}(h_{t}|\textbf{{h}}_{<t},\mathcal{S},\theta).$ (2)
+Here $\theta$ is the model parameter, $d$ refers to the task description and $w_{t}$ denotes the weight of the token at position $t$, where previous works normally treat each token equally Yuan et al.
 (2021); Fu et al.
-(2023); Finally, $\mathcal{S}$ is determined according to the aspect to evaluate.
+(2023).
+We provide further descriptions of each type of evaluator below.
 BARTScore
@@
 (2023), showing a great performance and covering a rich variety of aspects for comprehensive evaluations.
-With an understanding of natural language instructions, GPTScore (including GPT-X and FLAN-T5 models) can perform intricate and personalised assessments without requiring additional training.
+With an understanding of natural language instructions, GPTScore (including GPT-X and FLAN-T5 models) can perform intricate and personalized assessments without additional training.
 Cohere
-We additionally include Cohere, the more recent language model to enrich our experiment.
+We additionally include Cohere, the more recent language model to enrich our assessments.
 The evaluation scores assigned by the model is calculated according to Eq.
@@
 To ensure the reproducibility of our analysis, we exclusively employ publicly available checkpoints for the utilized models.
-Apart from the GPT3-Curie model that is taken from the OpenAI API [^3] Note that the newer models offered by the OpenAI API, such as GPT-4 , no longer support providing token probabilities, and hence they cannot be tested in our experiment.
-and generation model obtained from Cohere, the rest of the models are taken from the Hugging Face model hub [^4] https://huggingface.co/models..
+Apart from the GPT3-Curie model that is taken from the OpenAI API and generation model obtained from Cohere, the rest of the models are taken from the Hugging Face model hub [^3] https://huggingface.co/models..
 We use each of these generation models both for generating the summaries [^4] We use the zero-shot setting for the models that are not finetuned on summarization datasets.
@@
 These benchmarks include summaries from various generation models, as well as human evaluations, enabling us to assess the quality of these summaries.
-The SummEval benchmark contains summaries generated by various summarization models for 100 articles from the CNN/DM test set, with each summary supplemented by human annotations.
+The SummEval benchmark contains summaries generated by various summarization models (i.e.
+BART, T5 and GPT2) for 100 articles from the CNN/DM test set, with each summary supplemented by human annotations.
 More specifically, SummEval incorporates human annotations by both expert and crowd-sourced human annotators, targeting dimensions of coherence, consistency, fluency, and relevance.
@@
 We use 100 summaries from each of the BART and GPT-3 models from the ROSE benchmark.
-The RoSE benchmark proposed an assessment protocol termed ”Atomic Content Units” (ACUs).
+The RoSE benchmark proposed an assessment protocol termed “Atomic Content Units” (ACUs) Liu et al.
+(2022).
 ACU score gauges quality of evaluated summaries based on whether the presence of single facts (i.e., atomic facts) from reference are included in the evaluated summaries.
@@
 ${f(s,\mathcal{A})=\frac{\lvert\mathcal{A}_{s}\rvert}{\lvert\mathcal{A}\rvert}}$ (3)
-where $\mathcal{A}$ is a set of ACUs from gold summaries and s is the candidate summary.
+where $\mathcal{A}$ is a set of ACUs from gold summaries and $\mathcal{A}_{s}$ denotes the ACUs of candidate summary s.
 Max Min Mean Median
@@
 RoSE-GPT3 0.90 0.00 0.27 0.25
-SummEval-BART 5.00 2.67 4.56 4.67
-SummEval-T5 5.00 2.33 4.46 4.67
-SummEval-GPT2 5.00 1.33 3.52 3.33
-Table 1: Distribution of human annotation scores on RoSE and SummEval datasets, where RoSE mainly considers the index ’ACU’ and SummEval focuses on ’Coherence’ and ’Consistency’ aspects according to expert annotators.
+SummEval-BART 5.00 2.67 4.57 4.67
+SummEval-T5 5.00 2.33 4.52 4.67
+SummEval-GPT2 5.00 1.33 3.57 3.58
+Table 1: Distribution of human annotation scores on the RoSE and SummEval datasets, where in RoSE we consider the ‘ACU’ score, and in SummEval we focus on four aspects—‘Coherence’, ‘Consistency’, ‘Fluency’, and ‘Relevance’—as evaluated by expert annotators.
+The scores for SummEval are obtained by averaging the scores across all aspects and evaluations from all annotators.
 The distribution of human scores in RoSE and SummEval are given in Table 1.
@@
 All evaluator scores are normalized to a range between 0 and 1.
-NB: the number in the bracket represents the average length of summaries (measured in words) produced by the respective model.
+Additionally, the number in the bracket represents the average length of summaries (measured in words) produced by the respective model.
 3.4 Quantitative Comparisons
@@
 We assess the evaluators in two settings: (a) reference-free, where the metric evaluates the likelihood of the summary being generated from the source text, and (b) reference-based, where the generated summary is evaluated based on the reference summary.
-Due to the nature of log probabilities, original scores from each evaluator should be negative, and a higher score indicates better quality according to the evaluator.
+Due to the nature of log probabilities, original scores from each evaluator is be negative, and a higher score indicates better quality according to the evaluator.
 When weights $w_{t}$ in Eq.
@@
 as formulated in Eq.
-4, where $S_{i,j}$ indicates scores evaluated by the $j$-th evaluator on summaries generated by the $i$-th generator.
-$S_{i,j}^{norm}=\frac{\alpha(S_{i,j}-\min_{i}S_{i,j})}{\max_{i}S_{i,j}-\min_{i}S_{i,j}}$ (4)
-In this context, a normalized score of $\alpha$ signifies the highest quality attributed by the generator, while a score of 0 indicates the lowest quality.
+4, where $X_{i,j}$ indicates scores evaluated by the $j$-th evaluator on summaries generated by the $i$-th generator.
+$X_{i,j}^{norm}=\frac{\alpha(X_{i,j}-\min_{i}X_{i,j})}{\max_{i}X_{i,j}-\min_{i}X_{i,j}}$ (4)
+In this context, a normalized score of $\alpha$ signifies the highest quality attributed by the evaluator, while a score of 0 indicates the lowest quality.
+As the length of the generated summary is a key factor influencing the evaluation results, we further analyse the impact of lengths for the content generated by the models along with the experiments.
+In this regard, we also compute the correlations between the length of the text and the scores assigned by evaluators to identify trends in evaluators’ preferences.
 3.5 Qualitative Analysis
@@
 (2007), which respectively assess monotonic relationships and order associations between human evaluations and LM evaluator scores.
-For the SummEval dataset, we calculate the average value of expert annotations for coherence and consistency aspects, aligned with the reference-free input setting in the evaluation protocol specified in Yuan et al.
+They are common metrics for assessing correlations with human judgements.
+For the SummEval dataset, we calculate the correlations for four aspects (i.e.
+Coherence, Consistency, Fluency and Relevance ), aligned with the reference-free input setting in the evaluation protocol as specified by Yuan et al.
 (2021).
-For the evaluations based on the RoSE benchmark, ACU annotations are suited for reference-based summary salience evaluation.
+For the evaluations based on the RoSE benchmark, we use ACU annotations that are suited for reference-based summary salience evaluation.
 Therefore, we employ the correlation values obtained from the SummEval dataset for the reference-free setting and those from the RoSE benchmark for the reference-based setting.
 4 Experimental Results
-4.1 Quantitative Comparisons: Assessing Bias in LM-Evaluators Towards Their Own Output
+4.1 Quantitative Comparisons: Assessing Self-Bias in LM-Evaluators Towards Their Own Output
 Figure 3: Assessing Bias on CNN/DM Dataset using heatmaps in the reference-based setting.
-Observing darker cells along the diagonal line indicates potential bias.
+Observing darker cells along the diagonal line indicates potential self-bias.
 All evaluator scores are normalized to a range between 0 and 1.
-NB: the number in the bracket represents the average length of summaries (measured in words) produced by the respective model.
-Figures 2 and 3 display heatmaps presenting evaluator scores for various summaries generated by different generators from CNN/DM documents.
+Additionally, the number in the bracket represents the average length of summaries (measured in words) produced by the respective model.
+Figures 2 and 3 display heatmaps presenting evaluator scores for various summaries generated by different generators from CNN/DM documents in reference-free and reference-base settings, respectively.
 These scores are computed by averaging the individual scores of the selected 500 documents.
 In both heatmaps, we observe darker cells along the diagonal line, running from the top left to the bottom right.
-This indicates potential evaluator bias towards their corresponding generator models.
+This indicates the potential evaluator bias towards their corresponding generator models i.e., self-bias.
 However, this bias is notably more pronounced in the reference-free setting, commonly used for factuality evaluation Koh et al.
 (2022).
-In this context, we highlight the presence of the darkest cells along the diagonal line.
 Furthermore, as shown in Figure 2, we note a distinct trend: T5-based generators, whether fine-tuned or not, tend to receive higher scores when assessed using different T5Score variations compared to evaluations using BARTScore, GPTScore, or Cohere.
@@
 Similarly, we observe a parallel trend for BART-based generators, whether fine-tuned or not.
-It is also quite evident that when using FLAN-T5 variants and Cohere as evaluators, they tend to favor the summaries generated by their respective models.
-Meanwhile, evaluators are prone to assigning higher ranks to generators under the same dataset compared to those fine-tuned on others.
-For instance, when utilizing T5 fine-tuned on the XSUM dataset as evaluators, a preference is observed for BART-XSUM generators over T5-vanilla models.
+Meanwhile, evaluators tend to assign higher ranks to generators trained on the same dataset as themselves, rather than to those fine-tuned on different datasets (see Figure 6).
+For example, when using T5 models fine-tuned on the XSUM dataset as evaluators, there is a noticeable preference for BART-XSUM generators over T5-vanilla models, even though the evaluations are performed for the CNN Daily dataset.
 We observe the same pattern on summaries generated based XSUM documents.
-The corresponding heatmap is included in Figure 6 in Appendix B.1.
-It can also be observed that BART-based evaluators tend to rank FLAN-T5 generators the lowest, resulting in a normalized score of 0.
-This is likely to be attributed to the difference between BART and FLAN-T5’s architectures and BART-based models’ preference towards longer summaries.
-Similar observations can be observed that T5-based models score the summaries generated by Cohere low, as they tend to favor shorter summaries.
-Detailed analysis of the correlation between the length of generated summaries and the scores assigned by each evaluator is given in Figure 5.
-Figure 4: Heatmaps of evaluation scores on the RoSE & SummEval benchmarks for the reference-free and reference-based setting.
-It is recommended to use the reference-free setting for SummEval and the reference-based setting for RoSE, aligning with the specific aspects each benchmark emphasizes.
+Figure 4: Heatmaps of evaluation scores on the SummEval & RoSE benchmarks for the reference-free and reference-based setting.
+We use the reference-free setting for SummEval and the reference-based setting for RoSE, aligning with the specific aspects each benchmark emphasizes.
 Figure 5: Spearman Correlation between the length of generated summaries and the reference-free scores assigned by each evaluator.
 A higher positive score indicates that an evaluator prefers longer summaries, while a lower negative score indicates a preference for shorter summaries.
+Figure 6: Assessing Bias in the XSUM Dataset using heatmaps in the reference-free setting.
+Observing the darkest cells along the diagonal line, from the top left to the bottom right, indicates a distinct bias among evaluators towards their respective models.
+All evaluator scores are normalized to a range between 0 and 1.
+Additionally, the number in the bracket represents the average length of summaries (measured in words) produced by the respective model.
 4.2 Bias towards Longer Summaries
@@
 To further investigate this phenomenon, we calculate the average length of summaries generated by each generator for each of the datasets.
-Notably, BART models and the Cohere that have not been fine-tuned for summarization tend to produce the longest summaries on average.
+Notably, BART models and Cohere that have not been fine-tuned for summarization tend to produce the longest summaries on average.
 This is followed by the fine-tuned BART models on the CNN dataset.
-Conversely, models fine-tuned on XSUM generate the shortest summaries.
-[^8] Note that the XSUM dataset contains one-sentence summaries.
+Conversely, T5-based models score the summaries generated by Cohere low, as they tend to favour shorter summaries.
+A similar preference for short summaries can also be observed for evaluators fine-tuned on XSUM, which one-sentence summaries.
 Subsequently, we computed the Spearman correlation between the scores under the reference-free setting given by each of our examined evaluators and the length of the corresponding summary.
@@
 We observe a similar but less obvious pattern in the reference-based evaluations, as shown in Figure 3.
-4.3 Qualitative Analysis: Correlation of Bias with Human Evaluation
-To further verify the same model bias, we repeat the experiments from § 4.1 on summarization benchmarks that are accompanied by human evaluations.
+4.3 Qualitative Analysis: Correlation of Self-Bias with Human Evaluation
+To further verify the evaluators’ self-bias, we repeat the experiments from § 4.1 on summarization benchmarks that are accompanied by human evaluations.
 While the number of summaries in these benchmarks is limited compared to those in § 4.1, we can use the human annotations to verify that the inflated scores are not correlated with human evaluations.
-Figure 4 shows the evaluation results for the SummEval and RoSE benchmarks for the reference-free and reference-based setting respectively.
-It is advisable to utilize the reference-free setting for SummEval and the reference-based setting for RoSE, aligning with the specific aspects each benchmark emphasizes Yuan et al.
+Figure 4 shows the evaluation results for the SummEval and RoSE benchmarks for the reference-free and reference-based setting, respectively.
+As mentioned, we use SummEval for the reference-free setting and RoSE for the reference-based setting with regard to the specific aspects of each of these benchmarks Yuan et al.
 (2021).
@@
 Meanwhile, BART-based models receive higher scores from both BARTScore and GPTScore evaluators, instead of T5 evaluator.
-For human correlations, we use the RoSE and SummEval benchmark for evaluating the correlation in the reference-based and reference-free setting respectively.
-Table 3 (in Appendix) presents the Spearman and Kendall correlation values of SummEval in the reference-free setting.
+Table 3 presents the Spearman and Kendall correlation values of SummEval in the reference-free setting, whereas the Spearman and Kendall correlation values of RoSE in the reference-based setting are given in Table 2.
 Overall, we observe that none of the evaluators have a strong correlation with the human annotations on either of these benchmarks.
@@
 Nevertheless, these results demonstrate that none of these evaluators highly correlate with human annotations, and as observed in § 4.1, their inflated scores for their own underlying generator may contribute to this low correlation.
-Evaluator RoSE - Reference-based
+RoSE - Reference-based
+Evaluator ACU
 Spearman Kendall
@@
 Cohere-Command 0.384 0.267
-Table 2: Spearman and Kendall correlations between reference-based evaluation scores and human annotations using annotations in RoSE.
-The highest coefficient in each human annotation setting is bolded.
+Table 2: Spearman and Kendall correlations between reference-based evaluation scores and human annotations using annotations in RoSE. Results in bold indicate the strongest coefficient.
+SummEval - Reference-free
+Evaluator Coherence Consistency Fluency Relevance
+Spearman Kendall Spearman Kendall Spearman Kendall Spearman Kendall
+BART-Base -0.028 -0.021 0.107 0.078 -0.043 -0.037 0.105 0.074
+BART-Large 0.052 0.040 0.180 0.137 0.053 0.037 0.180 0.128
+BART-Base-CNN 0.193 0.138 0.228 0.171 0.190 0.145 0.069 0.050
+BART-Large-CNN 0.171 0.119 0.255 0.192 0.156 0.119 0.157 0.111
+BART-Base-XSUM 0.170 0.120 -0.103 -0.079 0.068 0.055 -0.174 -0.124
+BART-Large-XSUM 0.055 0.040 0.060 0.046 -0.025 -0.022 0.080 0.056
+T5-Small 0.208 0.146 0.547 0.419 0.501 0.398 0.415 0.295
+T5-Base 0.173 0.119 0.533 0.409 0.488 0.381 0.367 0.260
+T5-Large 0.185 0.132 0.477 0.364 0.445 0.345 0.387 0.281
+T5-Small-CNN 0.315 0.222 0.462 0.356 0.401 0.314 0.299 0.214
+T5-Base-CNN 0.192 0.135 0.253 0.190 0.189 0.150 0.148 0.106
+T5-Small-XSUM 0.245 0.178 0.142 0.109 0.209 0.164 0.113 0.079
+T5-Large-XSUM 0.213 0.152 -0.111 -0.085 0.018 0.012 -0.041 -0.029
+GPT2 0.103 0.077 0.154 0.117 0.037 0.026 0.032 0.021
+GPT2-Medium 0.123 0.091 0.234 0.179 0.117 0.086 0.066 0.047
+GPT2-Large 0.119 0.089 0.184 0.140 0.107 0.080 0.024 0.017
+GPT3-Curie 0.152 0.108 0.483 0.371 0.345 0.264 0.311 0.223
+FLANT5-Base 0.220 0.154 0.448 0.345 0.295 0.228 0.229 0.159
+FLANT5-XL 0.248 0.174 0.550 0.424 0.389 0.301 0.402 0.289
+Cohere-Command 0.136 0.097 0.520 0.397 0.351 0.268 0.427 0.302
+Table 3: Spearman and Kendall correlations between the reference-free evaluation scores and expert annotations provided in SummEval on four different aspects.
+The strongest correlation for each aspect is bolded.
 5 Conclusions
-Based on experiments, we make the following conclusions:
-First, the popularity of generative evaluation metrics, such as BARTScore, is on the rise for evaluating the factual accuracy of generated content—a critical concern in modern generator models.
-However, our results reveal that this evaluation approach is susceptible to the same model bias, highlighting the need for more robust metrics to assess factual correctness reliably.
+Based on experiments, we make the following conclusions: First, the popularity of generative evaluation metrics, such as BARTScore, is on the rise for evaluating the factual accuracy of generated content—a critical concern in modern generator models.
+However, our results reveal that this evaluation approach is susceptible to the self-bias, highlighting the need for more robust metrics to assess factual correctness reliably.
 Second, our analysis indicates that models fine-tuned on the XSUM dataset are not suitable for direct integration into evaluators due to their bias towards shorter summaries.
@@
 This bias should be considered when interpreting and applying these metrics.
-Finally, our study uncovers the presence of the same model bias across all assessed evaluators.
+Finally, our study uncovers the presence of the self-bias across all assessed evaluators.
 Consequently, we recommend avoiding the use of the same underlying model as the generator for assessment.
@@
 Limitations
-We noted that our work has the following limitations.
+We note that our work has the following limitations.
 Firstly, our experiment has been focused on the summarization task.
@@
 Ethics Statement
-This paper raises no ethical concerns. The data and additional resources used in this study are open-source and widely employed in existing works.
+This paper raises no ethical concerns. The data and supplementary materials used in this study are open-sourced and widely employed in existing works.
 Appendix A Evaluation Setting
-Evaluator SummEval - Reference-free
-Spearman Kendall
-BART-Base 0.023 0.010
-BART-Large 0.120 0.082
-BART-Base-CNN 0.240 0.164
-BART-Large-CNN 0.227 0.149
-BART-Base-XSUM 0.081 0.055
-BART-Large-XSUM 0.061 0.043
-T5-Small 0.396 0.275
-T5-Base 0.365 0.249
-T5-Large 0.337 0.236
-T5-Small-CNN 0.448 0.315
-T5-Base-CNN 0.259 0.172
-T5-Small-XSUM 0.235 0.167
-T5-Large-XSUM 0.110 0.081
-GPT2 0.153 0.107
-GPT2-Medium 0.207 0.146
-GPT2-Large 0.183 0.128
-GPT3-Curie 0.326 0.228
-FLANT5-Base 0.361 0.252
-FLANT5-XL 0.433 0.305
-Cohere-Command 0.328 0.227
-Table 3: Spearman and Kendall correlations between reference-free evaluation scores and human annotations using annotations in SummEval.
-The highest coefficient in each human annotation setting is bolded.
 A.1 Generator
-Further information about models utilized in our generation models including checkpoint names, prompt setting, etc., are presented in Table 4.
+Full details of the models (e.g. checkpoint, prompt setting) that we employed as generators are given in Table 4.
 Name of Generator Name of Checkpoint or Model Suffix Prefix
@@
 Cohere-Command api.cohere.ai/v1/generate Write a concise summarization:
-Table 4: Name of checkpoints or model utilized in our generation setting with corresponding prompt configurations, ’text-curie-001’ is the model name provided by OpenAI API, and ’api.cohere.ai/v1/generate’ denotes model names provided by Cohere API, alongside other checkpoints available through Hugging Face.
+Table 4: Checkpoints or model utilized in our generation setting with corresponding prompt configurations, ‘text-curie-001’ is the model name provided by OpenAI API, and ‘api.cohere.ai/v1/generate’ denotes model names provided by Cohere API, alongside other checkpoints available through Hugging Face.
 A.2 Evaluator
-Further information about models utilized in our evaluation models are presented in Table 5.
+Full details of the models that we employed as our evaluators are given in Table 5 (reference-free settings) and Table 6 (reference-based settings).
 Name of Evaluator Name of Checkpoint or Model Suffix Prefix
@@
 Cohere-Command api.cohere.ai/v1/generate Write a concise summarization:
-Table 5: Name of checkpoints or model utilized in our evaluation study with corresponding prompt configurations, ’text-curie-001’ is the model name provided by OpenAI API, and ’api.cohere.ai/v1/generate’ denotes model names provided by Cohere API, alongside other checkpoints available through Hugging Face.
+Table 5: Checkpoints or model utilized in our evaluation study for the reference-free setting with corresponding prompt configurations, ‘text-curie-001’ is the model name provided by OpenAI API, and ‘api.cohere.ai/v1/generate’ denotes model names provided by Cohere API, alongside other checkpoints available through Hugging Face.
+Name of Evaluator Name of Checkpoint or Model Suffix Prefix
+BART-Base facebook/bart-base in other words:
+BART-Large facebook/bart-large in other words:
+BART-Base-CNN ainize/bart-base-cnn
+BART-Large-CNN facebook/bart-large-cnn
+BART-Base-XSUM morenolq/bart-base-xsum
+BART-Large-XSUM facebook/bart-large-xsum
+T5-Small t5-small Paraphrase:
+T5-Base t5-base Paraphrase:
+T5-Large t5-large Paraphrase:
+T5-Small-CNN ubikpt/t5-small-finetuned-cnn
+T5-Base-CNN flax-community/t5-base-cnn-dm
+T5-Small-XSUM pki/t5-small-finetuned xsum
+T5-Large-XSUM sysresearch101/t5-large-finetuned-xsum
+GPT2 openai-community/gpt2 Paraphrase the sentence:
+GPT2-Medium openai-community/gpt2-medium Paraphrase the sentence:
+GPT2-Large openai-community/gpt2-large Paraphrase the sentence:
+GPT3-Curie text-curie-001 Paraphrase the sentence:
+FLANT5-Base google/flan-t5-base Paraphrase the sentence:
+FLANT5-XL google/flan-t5-xl Paraphrase the sentence:
+Cohere-Command api.cohere.ai/v1/generate Paraphrase the sentence:
+Table 6: Checkpoints and models utilised in our evaluation study for the reference-based setting with corresponding prompt configurations, ‘text-curie-001’ is the model name provided by OpenAI API, and ‘api.cohere.ai/v1/generate’ denotes model names provided by Cohere API, alongside other checkpoints available through Hugging Face.
 Appendix B Evaluation Results
@@
 For the meta evaluation, Spearman and Kendall correlation values in the reference-free setting for SummEval benchmark are shown in Table 3.
-Figure 6: Assessing Bias in the XSUM Dataset using heatmaps in the reference-free setting.
-Observing the darkest cells along the diagonal line, from the top left to the bottom right, indicates a distinct bias among evaluators towards their respective models.
-All evaluator scores are normalized to a range between 0 and 1.
-NB: the number in the bracket represents the average length of summaries (measured in words) produced by the respective model.
 B.2 Reference-based Setting